# PokéLLMon: A Human-Parity Agent for Pokémon Battles with Large Language Models

### v2

@@ Abstract
We introduce PokéLLMon, the first LLM-embodied agent that achieves human-parity performance in tactical battle games, as demonstrated in Pokémon battles. The design of PokéLLMon incorporates three key strategies: (i) In-context reinforcement learning that instantly consumes text-based feedback derived from battles to iteratively refine the policy; (ii) Knowledge-augmented generation that retrieves external knowledge to counteract hallucination and enables the agent to act timely and properly; (iii) Consistent action generation to mitigate the panic switching phenomenon when the agent faces a powerful opponent and wants to elude the battle. We show that online battles against human demonstrates PokéLLMon’s human-like battle strategies and just-in-time decision making, achieving 49% of win rate in the Ladder competitions and 56% of win rate in the invited battles. Our implementation and playable battle logs are available at: https://github.com/git-disl/PokeLLMon.
@@ 1 Introduction
Games are suitable test-beds to develop LLM-embodied agents (Duan et al., 2022; Batra et al., 2020) to interact with the virtual environment in a way resembling human behavior. For example, Generative Agents (Park et al., 2023) conducts a social experiments with LLMs assuming various roles in a “The Sims”-like sandbox, where agents exhbit behavior and social interactions mirroring humans. In Mincraft, decision-making agents (Wang et al., 2023a, c; Singh et al., 2023) are designed to explore the world and develop new skills for solving tasks and making tools.
@@
Scope and Contributions: The scope of this paper is to develop an LLM-embodied agent that mimics the way a human player engages in Pokémon battles. The objective is to explore the key factors that make the LLM-embodied agent a good player and to examine its strengths and weaknesses in battles against human players. To enable LLMs play game autonomously, we implement an environment that can parse and translate battle state into text description, and deliver generated action back to the server. By evaluating existing LLMs, we identify the presence of hallucination, and the panic switching phenomenon.
@@
  • PokéLLMon, to the best of our knowledge, is the first LLM-embodied agent with human-parity performance in tactical battle games.
@@ 2 LLMs as Game Players
Open-ended games: Open-ended games allow players to freely explore the game world and interact with others. Generative Agent (Park et al., 2023) showcases that LLM-embodied agents exhibit behavior and social interactions mirroring human-like patterns. In MineCraft, Voyager (Wang et al., 2023a) employs curriculum mechanism to explore the world and generates and executes code for solving tasks. DEPS (Wang et al., 2023c) proposes an approach of “Describe, Explain, Plan and Select” to accomplish 70+ tasks. Planing-based frameworks like AutoGPT (Significant Gravitas, ) and MetaGPT (Hong et al., 2023) can be adopted for the exploration task as well.
@@ 5.1 Pokémon Battles
Player Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@ 6.1 In-Context Reinforcement Learning (ICRL)
Player Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@ 6.2 Knowledge-Augmented Generation (KAG)
Player Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@ 6.3 Consistent Action Generation
Player Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@
Player Win rate normal-$\uparrow$ Switch rate CS1 rate CS2 rate
@@ 7.1 Battle Against Human Players
v.s. Player Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@ 7.2 Battle Skill Analysis
Ladder Win rate normal-$\uparrow$ Score normal-$\uparrow$ Turn # Battle #
@@ 8 Conclusion
In this paper, we enable LLMs to autonomously play the well-known Pokémon battles against human. We introduce PokéLLMon, the first LLM-embodied agent that achieves human-competent performance in tactical battle games. We introduce three key strategies in the design of PokéLLMon: (i) In-Context Reinforcement Learning, which consumes the text-based feedback as “reward” to iteratively refine the action generation policy without training; (ii) Knowledge-Augmented Generation that retrieves external knowledge to combat hallucination and ensures the agent act timely and properly; (iii) Consistent Action Generation that prevents the panic switching issue when encountering powerful opponents. The architecture of PokéLLMon is general and can be adapted for the design of LLM-embodied agents in many other games, addressing the problems of hallucination and action inconsistency.

### v3

@@ Abstract
We introduce PokéLLMon, the first LLM-based agent that achieves human-parity performance in tactical battle games, as demonstrated in Pokémon battles. The design of PokéLLMon incorporates three key strategies: (i) In-context reinforcement learning that instantly consumes text-based feedback derived from battles to iteratively refine the policy; (ii) Knowledge-augmented generation that retrieves external knowledge to counteract hallucination and enables the agent to act timely and properly; (iii) Consistent action generation to mitigate the panic switching phenomenon when the agent faces a powerful opponent and wants to elude the battle. We show that online battles against human demonstrates PokéLLMon’s human-like battle strategies and just-in-time decision making, achieving 49% of win rate in the Ladder competitions and 56% of win rate in the invited battles. Our implementation and playable battle logs are available at: https://github.com/git-disl/PokeLLMon.
@@ 1 Introduction
Games are suitable test-beds to develop LLM-based agents (Duan et al., 2022; Batra et al., 2020) to interact with the virtual environment in a way resembling human behavior. For example, Generative Agents (Park et al., 2023) conducts a social experiments with LLMs assuming various roles in a “The Sims”-like sandbox, where agents exhbit behavior and social interactions mirroring humans. In Mincraft, decision-making agents (Wang et al., 2023a, c; Singh et al., 2023) are designed to explore the world and develop new skills for solving tasks and making tools.
@@
Scope and Contributions: The scope of this paper is to develop an LLM-based agent that mimics the way a human player engages in Pokémon battles. The objective is to explore the key factors that make the LLM-based agent a good player and to examine its strengths and weaknesses in battles against human players. To enable LLMs play game autonomously, we implement an environment that can parse and translate battle state into text description, and deliver generated action back to the server. By evaluating existing LLMs, we identify the presence of hallucination, and the panic switching phenomenon.
@@
  • PokéLLMon, to the best of our knowledge, is the first LLM-based agent with human-parity performance in tactical battle games.
@@ 2 LLMs as Game Players
Open-ended games: Open-ended games allow players to freely explore the game world and interact with others. Generative Agent (Park et al., 2023) showcases that LLM-based agents exhibit behavior and social interactions mirroring human-like patterns. In MineCraft, Voyager (Wang et al., 2023a) employs curriculum mechanism to explore the world and generates and executes code for solving tasks. DEPS (Wang et al., 2023c) proposes an approach of “Describe, Explain, Plan and Select” to accomplish 70+ tasks. Planing-based frameworks like AutoGPT (Significant Gravitas, ) and MetaGPT (Hong et al., 2023) can be adopted for the exploration task as well.
@@ 5.1 Pokémon Battles
Player Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@ 6.1 In-Context Reinforcement Learning (ICRL)
Player Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@ 6.2 Knowledge-Augmented Generation (KAG)
Player Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@ 6.3 Consistent Action Generation
Player Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@
Player Win rate $\uparrow$ Switch rate CS1 rate CS2 rate
@@ 7.1 Battle Against Human Players
v.s. Player Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@ 7.2 Battle Skill Analysis
Ladder Win rate $\uparrow$ Score $\uparrow$ Turn # Battle #
@@ 8 Conclusion
In this paper, we enable LLMs to autonomously play the well-known Pokémon battles against human. We introduce PokéLLMon, the first LLM-based agent that achieves human-competent performance in tactical battle games. We introduce three key strategies in the design of PokéLLMon: (i) In-Context Reinforcement Learning, which consumes the text-based feedback as “reward” to iteratively refine the action generation policy without training; (ii) Knowledge-Augmented Generation that retrieves external knowledge to combat hallucination and ensures the agent act timely and properly; (iii) Consistent Action Generation that prevents the panic switching issue when encountering powerful opponents. The architecture of PokéLLMon is general and can be adapted for the design of LLM-based agents in many other games, addressing the problems of hallucination and action inconsistency.